OpenStudy (anonymous):

Give an example of a 2x2 matrix whose determinant is 13. How would I come up with numbers? Is it just a random guess and check?

jimthompson5910 (jim_thompson5910):

If you have the 2x2 matrix a b c d The determinant D is D = ad - bc

OpenStudy (anonymous):

How can I put that into actual numbers?

OpenStudy (anonymous):

1 0 0 a What is the deterninant?

OpenStudy (anonymous):

13

OpenStudy (anonymous):

1 0 0 13

OpenStudy (anonymous):

Nice lol. So is it just a guess and check?

OpenStudy (anonymous):

There are a lot of matrices that have determinant 13

jimthompson5910 (jim_thompson5910):

no, just fix 3 values and make the fourth value unknown and solve for it

OpenStudy (anonymous):

Yes an infinite number of matrices have determinant 13.

OpenStudy (anonymous):

What if you get a decimal?

OpenStudy (anonymous):

1 1 1 14

OpenStudy (anonymous):

How can you make sure that you don't get a decimal?

OpenStudy (anonymous):

1 2 1 15

OpenStudy (anonymous):

1 3 1 16

OpenStudy (anonymous):

1 4 1 17

OpenStudy (anonymous):

ok dude thanks, but how do you make sure you don't get a decimal? How are you coming up with so many

OpenStudy (anonymous):

If all your entries are integers, the determinant is an integer.

OpenStudy (anonymous):

1 17 1 30

jimthompson5910 (jim_thompson5910):

The determinant is D = ad - bc Fix values for a, b, and c. Say a = 2, b = 3, and c = 4 (these values are completely random) D = ad - bc 13 = 2d - 3*4 13 = 2d - 12 13+12 = 2d 25 = 2d 25/2 = d d = 12.5 So 2 3 4 12.5 has a determinant of 13

OpenStudy (anonymous):